Last year, the company sold 54,000 of these balls, with the following results:

Sales (54,000 balls) $ 1,350,000

Variable expenses 810,000

Contribution margin 540,000

Fixed expenses 372,000

Net operating income $ 168,000

Required: 1. Compute (a) last year's CM ratio and the break-even point in balls, and (b) the degree of operating leverage at last year’s sales level.

2. Due to an increase in labor rates, the company estimates that next year's variable expenses will increase by $3.00 per ball. If this change takes place and the selling price per ball remains constant at $25.00, what will be next year's CM ratio and the break-even point in balls?

3. Refer to the data in (2) above. If the expected change in variable expenses takes place, how many balls will have to be sold next year to earn the same net operating income, $168,000, as last year?

4. Refer again to the data in (2) above. The president feels that the company must raise the selling price of its basketballs. If Northwood Company wants to maintain the same CM ratio as last year (as computed in requirement 1a), what selling price per ball must it charge next year to cover the increased labor costs?

5. Refer to the original data. The company is discussing the construction of a new, automated manufacturing plant. The new plant would slash variable expenses per ball by 40.00%, but it would cause fixed expenses per year to double. If the new plant is built, what would be the company’s new CM ratio and new break-even point in balls?

6. Refer to the data in (5) above. a. If the new plant is built, how many balls will have to be sold next year to earn the same net operating income, $168,000, as last year? b. Assume the new plant is built and that next year the company manufactures and sells 54,000 balls (the same number as sold last year). Prepare a contribution format income statement and compute the degree of operating leverage.
Answer #1

1. Compute (a) last year's CM ratio and the break-even point in balls,

Variable expense per ball= $810000/54000= $15

Sales price per ball= $1350000/54000= $25

Contribution margin per ball= $25-15= $10

CM ratio= Contribution margin*100/Sales

= $10*100/25= 40%

Break even point in balls= Fixed cost/Contribution margin per ball

= $372000/10= 37200 balls

(b) the degree of operating leverage at last year’s sales level.

Degree of operating leverage= Contribution margin/Net operating income

= $540000/168000= 3.21

2. Due to an increase in labor rates, the company estimates that next year's variable expenses will increase by $3.00 per ball. If this change takes place and the selling price per ball remains constant at $25.00, what will be next year's CM ratio and the break-even point in balls?

New variable cost per ball= $15+3- $18

New contribution margin per ball= $25-18= $7

CM ratio= Contribution margin*100/Sales

= $7*100/25= 28%

Break even point in balls= Fixed cost/Contribution margin per ball

= $372000/7= 53143 balls

3. Refer to the data in (2) above. If the expected change in variable expenses takes place, how many balls will have to be sold next year to earn the same net operating income, $168,000, as last year?

Sales in balls= (Fixed cost+Target income)/Contribution margin per ball

= ($372000+168000)/7= 77143

4. Refer again to the data in (2) above. The president feels that the company must raise the selling price of its basketballs. If Northwood Company wants to maintain the same CM ratio as last year (as computed in requirement 1a), what selling price per ball must it charge next year to cover the increased labor costs?

Let the selling price per ball= X

Contribution margin per ball= X*40%= 0.4X

Variable cost= Selling price-Contribution margin= X-0.4X= 0.6X

Variable cost= $18

0.6X= $18

X= $30

Selling price per ball= X= $30

5. Refer to the original data. The company is discussing the construction of a new, automated manufacturing plant. The new plant would slash variable expenses per ball by 40.00%, but it would cause fixed expenses per year to double. If the new plant is built, what would be the company’s new CM ratio and new break-even point in balls?

New variable cost per ball= ($15*60%)= $9

New fixed costs= $372000*2= $744000

Contribution margin per ball= Selling price per ball-New variable cost per ball

= $25-9= $16

Contribution margin ratio= Contribution margin*100/Sales

= $16*100/25= 64%

Break even point in balls= Fixed cost/Contribution margin per unit

= $744000/16= 46500 balls

6. Refer to the data in (5) above. a. If the new plant is built, how many balls will have to be sold next year to earn the same net operating income, $168,000, as last year?

6-A) Break eve point in balls= (Fixed cost+Target income)/Contribution margin per unit

= ($744000+168000)/16= 57000 balls

b. Assume the new plant is built and that next year the company manufactures and sells 54,000 balls (the same number as sold last year). Prepare a contribution format income statement and compute the degree of operating leverage.

Northwood Company
Contribution Income Statment
Sales (54000*$25) $1350000
Variable expenses (54000*$9) -486000
Contribution margin 864000
Fixed expenses -744000
Net operating income $120000
Degree of operating leverage (864000/120000) 7.2